Solve by the addition method (elimination)
2x - 4y = 24
3x + 5y = 14

Found 2 solutions by hkwu, MathTherapy:
Answer by hkwu(60)

We have
3x + 5y + (2x -4y) = 24 + 14 (#2 + #1)
5x + y = 38
y = 38 - 5x
Substitute back into one of the original equations (#1 or #2, it doesn't matter; I will substitute into #2), to get
3x + 5y = 14
3x + 5(38 - 5x) = 14
3x + 190 - 25x = 14
190 - 22x = 14
22x = 176
x = 8
Now do one final substitution to get the value of y.
y = 38 - 5x
y = 38 - 5(8)
y = -2
Your solution is x = 8, y = -2.

Answer by MathTherapy(10552)


Solve by the addition method (elimination)
2x - 4y = 24
3x + 5y = 14
2x - 4y = 24 ------ eq (i)
3x + 5y = 14 ------ eq (ii)
Multiply eq (i) by - 3 and eq (ii) by 2 to get eqs (iii) & (iv)
Add eqs (iii) & (iv) to eliminate x and determine the value of y
Substitute value of y into any of the 2 ORIGINAL EQUATIONS to determine the value of x
Still having difficulties? Write back!!